# Real-time, Adaptive Radiological Anomaly Detection and Isotope Identification Using Non-negative Matrix Factorization

### v1

@@ I Introduction
Non-negative matrix factorization (NMF) [50] has been shown to be a useful framework for modeling gamma-ray background signatures [27, 17, 51, 52] when using “global” models trained on large datasets containing data representative of background signatures present in test cases. However, the existing Berkeley Lab-developed NMF-based approach [51] lacks the flexibility required for deployment on mobile detector systems in scenarios where a pre-trained model is unable to capture the complexity of unseen data. In Section II, we will discuss why NMF-based algorithms are inherently well suited for detection and identification tasks and how the constraints of the conventional approach lead to clear limitations of performance on mobile detection tasks. We then present Adaptive NMF in Section III, an update to the conventional approach that seeks to address the issue of temporal drift and extend the NMF framework to unseen environments. We then show in Section IV that Adaptive NMF achieves anomaly detection performance superior to other algorithms commonly used in the field, while maintaining a low FAR under rapidly changing environmental conditions.
@@ II-A Background Modeling with NMF
where is the dimension of each of the vectors, is the number of vectors in the dataset, and is the rank of the resulting approximation matrix, . The columns of the matrix V can be interpreted as a set of basis vectors for a lower-dimensional space that contains the Poisson expectation values of the original training data points111The space is not a true vector space because of the non-negative constraints. In particular, the set more closely resembles a conical hull with the additional constraints that all vectors in the set be non-negative.. The rows of A function as the coefficients for linear combinations of the basis vectors that approximate the members of the original dataset. By decomposing a set of gamma-ray spectra in this way, the columns of V can be interpreted as the constituent components of the local gamma-ray background signature. In order to force the NMF components to be consistent with Poisson statistics, the background model is fit by optimizing the Poisson negative log likelihood loss function:
@@ IV Results
Adaptive NMF was tested against gross counts k-sigma [37, 38], multiplexed censored energy window (mCEW) [39], conventional NMF [51], and Nuisance-Rejection Spectral Comparison Ratio (NSCRAD) [26, 35]. The implementations of the benchmark algorithms come from the RADAI algorithm library [53].
@@ IV-A Simulated Datasets
RADAI provides a dataset [57, 58] that simulates a 2”4”16” NaI detector moving through an urban environment. The purpose of the RADAI dataset is to provide users with simulated source encounters in environments where data collection would be prohibitively expensive or otherwise infeasible. It includes sources one might typically encounter in an urban environment like those used in medical and industrial applications, threat sources like highly enriched Uranium and weapons-grade Plutonium, and NORM. All sources also have varying shielding configurations common for their individual applications. Single-source encounters with all 21 of the sources listed in Table I are included in the dataset. The data were prepared by integrating list-mode events into a series of one-second spectra in 129 nonlinear energy bins from 45–3000 keV. The width of these bins is proportional to the square root of the energy in order to ensure that the ratio of the bin widths to the detector energy resolution is approximately constant throughout the entire spectrum. The Adaptive NMF template library was also generated from RADAI source data. For these tests, Adaptive NMF was instantiated with a background buffer length of 10000 1-second integrated spectra; a refit period of 600 seconds (10 minutes); integration times of 1, 2, 3, 4, and 5 seconds running in parallel; and a Tikhonov regularization factor of . The Conventional NMF instance was trained on 10 hours’ worth of background data (36000 spectra).
Fig. 5 shows a set of Receiver Operating Characteristic (ROC) curves comparing the detection performance of Adaptive NMF against the suite of RADAI benchmark algorithms tested on all 21 simulated sources. The x-axis of the ROC curve is the PFA per second, and the y-axis gives the probability of detection (PD) per source encounter. At functionally relevant false alarm rates (e.g., 1 per 8 hours, corresponding to a false alarm probability of about ), Adaptive NMF outperforms all benchmark methods.
@@ IV-B Field Tests
The following results come from tests performed on data collected in urban and rural locations with vehicle-borne 2”x4”x16” NaI detectors. We performed these tests with an instance of Adaptive NMF equipped with a background buffer of 3600 1-second integrated spectra; a refit period of 600 seconds (10 minutes); integration times of 1, 2, 4, and 8 seconds running in parallel; and a Tikhonov regularization factor of . Adaptive NMF was tested against Conventional NMF (trained on a dataset of 8742 spectra representative of the testing environment) and an implementation of NSCRAD that has been used in field applications. These tests consisted of encounters with four different sources at six different distances. Eight passes were performed at each distance.
@@ IV-C Nuisance Alarm Suppression
Adaptive NMF has a clear advantage over the other benchmark methods — particularly Conventional NMF — during rain events. As it begins to rain, there is a gradual increase in radon progeny in the background [59]. If a Conventional NMF instance faced such a scenario without having been previously trained on spectra collected during rain, it would likely raise false alarms as the concentration of radon exceeds the expected level. The false alarms would continue for hours after the rain stops until the radionuclides from the rain decay to a point where they no longer contribute to a significant component of the background.
@@ V Discussion
Perhaps unsurprisingly, the performance of the methods tested against the RADAI dataset, as shown in Fig. 5, follows the same hierarchy as the amount of information being leveraged in each method. K-sigma uses only gross count rate information and had the poorest performance, which was to be expected in the presence of a rapidly changing background. mCEW, which constructs spectral regions of interest from its training data to perform detection, had greater sensitivity than k-sigma but still suffered when faced with the highly variable RADAI background. NSCRAD, too, constructs spectral windows to detect broad classes of anomalies, but it also incorporates information about its environment via a continually updating background model. This enables it to outperform mCEW and k-sigma, but it is susceptible to becoming stuck in a state of false alarm when abrupt background shifts occur. Conventional NMF achieved greater sensitivity by preserving the most spectral information compared to the other methods described and by making use of a broad library of templates of potential sources and a background model trained on many hours of data. The fact that Adaptive NMF was able to, in most test cases, match or improve upon the detection sensitivity of Conventional NMF is particularly encouraging due to the tradeoff in the information being utilized by each method. Adaptive NMF builds its background models from much smaller datasets — on the order of 1 hour’s worth of data — while Conventional NMF commonly uses many hours of data to build a complex model that is capable of capturing the richness of an environment over time in a single model. We have shown that in the case of highly variable background compositions, a flexible local model is not only sufficient but oftentimes superior to a highly complex static model.

### v2

@@ I Introduction
Non-negative matrix factorization (NMF) [50] has been shown to be a useful framework for modeling gamma-ray background signatures [27, 17, 51, 52] when using “global” models trained on large datasets containing data representative of background signatures present in test cases. However, the existing Berkeley Lab-developed NMF-based approach [51] lacks the flexibility required for deployment on mobile detector systems in scenarios where a pre-trained model is unable to capture the complexity of unseen data. In Section II, we will discuss why NMF-based algorithms are inherently well suited for detection and identification tasks and how the constraints of the conventional approach lead to clear limitations of performance on mobile detection tasks. We then present Adaptive NMF in Section III, an update to the conventional approach that seeks to address the issue of temporal drift and extend the NMF framework to unseen environments.
Finally, Section IV we test the detection performance of Adaptive NMF against a suite of baseline algorithms including gross counts k-sigma [37, 38], multiplexed censored energy window (mCEW) [39], the ”conventional” NMF approach [51], and Nuisance-Rejection Spectral Comparison Ratio (NSCRAD) [26, 35]. 111These algorithms were selected as benchmark comparisons because they are commonly fielded in real-world applications, there exist open-source implementations, and were recommended by the Detecting Radiation Algorithms Group [53]. This is by no means an exhaustive list of algorithms used for anomaly detection tasks, and there exist methods used by entities in private industry. However, due to the proprietary nature of such methods, we were unable to test against them. We show that Adaptive NMF achieves anomaly detection performance superior to the benchmarks while maintaining a low FAR under rapidly changing environmental conditions.
@@ II-A Background Modeling with NMF
where is the dimension of each of the vectors, is the number of vectors in the dataset, and is the rank of the resulting approximation matrix, . The columns of the matrix V are a set of basis vectors for a lower-dimensional space that contains the Poisson expectation values of the original training data points222The space is not a true vector space because of the non-negative constraints. In particular, the set more closely resembles a conical hull with the additional constraints that all vectors in the set be non-negative.. In practice, the basis vectors can be interpreted as background spectral templates developed from the training process. The rows of A function as the coefficients for linear combinations of the basis vectors that approximate the members of the original dataset. By decomposing a set of gamma-ray spectra in this way, the columns of V can be interpreted as the constituent components of the local gamma-ray background signature. In order to force the NMF components to be consistent with Poisson statistics, the background model is fit by optimizing the Poisson negative log likelihood loss function:
@@ IV Results
Adaptive NMF was tested against gross counts k-sigma [37, 38], mCEW [39], Conventional NMF [51], and NSCRAD [26, 35]. The implementations of the benchmark algorithms come from the RADAI algorithm library [54].
@@ IV-A Simulated Datasets
RADAI provides a dataset [58, 59] that simulates a 2”4”16” NaI detector moving through an urban environment. The purpose of the RADAI dataset is to provide users with simulated source encounters in environments where data collection would be prohibitively expensive or otherwise infeasible. It includes sources one might typically encounter in an urban environment like those used in medical and industrial applications, threat sources like highly enriched Uranium and weapons-grade Plutonium, and NORM. All sources also have varying shielding configurations common for their individual applications. Single-source encounters with all 21 of the sources listed in Table I are included in the dataset. The data were prepared by integrating list-mode events into a series of one-second spectra in 129 nonlinear energy bins from 45–3000 keV. The width of these bins is proportional to the square root of the energy in order to ensure that the ratio of the bin widths to the detector energy resolution is approximately constant throughout the entire spectrum. The template library used by each of the benchmark algorithms that make use of spectroscopic information (i.e., all but k-sigma) was also generated from RADAI source data.
For these tests, Adaptive NMF was instantiated with a background buffer length of 10000 1-second integrated spectra; a refit period of 600 seconds (10 minutes); integration times of 1, 2, 3, 4, and 5 seconds running in parallel; and a Tikhonov regularization factor of . The Conventional NMF instance was trained on 10 hours’ worth of background data (36000 spectra).
Fig. 5 shows a set of Receiver Operating Characteristic (ROC) curves comparing the detection performance of Adaptive NMF against the suite of RADAI benchmark algorithms tested on all 21 simulated sources. The x-axis of the ROC curve is the PFA per second, and the y-axis gives the probability of detection (PD) per source encounter. Points on the curve are calculated by varying each algorithm’s detection threshold value and observing the corresponding PD and PFA. All other algorithm hyperparameter values are fixed. At functionally relevant false alarm rates (e.g., 1 per 8 hours, corresponding to a false alarm probability of about ), Adaptive NMF outperforms all benchmark methods.
@@ IV-B Field Tests
The following results come from tests performed on data collected in urban and rural locations with vehicle-borne 2”4”16” NaI detectors. We performed these tests with an instance of Adaptive NMF equipped with a background buffer of 3600 1-second integrated spectra; a refit period of 600 seconds (10 minutes); integration times of 1, 2, 4, and 8 seconds running in parallel; and a Tikhonov regularization factor of . Adaptive NMF was tested against Conventional NMF (trained on a dataset of 8742 spectra representative of the testing environment) and an implementation of NSCRAD that has been used in field applications. These tests consisted of encounters with four different sources at six different distances. Eight passes were performed at each distance.
@@ IV-C Nuisance Alarm Suppression
The Adaptive NMF algorithm is able to effectively adapt to changes in background environment and maintain a low and consistent false alarm rate compared to the conventional approach, as illustrated in Fig. 8. Both methods were trained on a dataset of spectra collected in Washington D.C. and then tested on a dataset of spectra collected in rural New Mexico. Conventional NMF (bottom) raises numerous false alarms due to the inconsistency of its background model with the environment on which it is being tested. In contrast, Adaptive NMF (top) begins analysis with a high false alarm rate as it gathers new background data, but is able to effectively capture the spectral signature in its background model after the first model refit.
Adaptive NMF has a clear advantage over the other benchmark methods — particularly Conventional NMF — during rain events. As it begins to rain, there is a gradual increase in radon progeny in the background [60]. If a Conventional NMF instance faced such a scenario without having been previously trained on spectra collected during rain, it would likely raise false alarms as the concentration of radon exceeds the expected level. Without operator intervention, the false alarms would continue for hours after the rain stops until the radionuclides from the rain decay to a point where they no longer contribute to a significant component of the background.
@@ V Discussion
Perhaps unsurprisingly, the performance of the methods tested against the RADAI dataset, as shown in Fig. 5, follows the same hierarchy as the amount of information being leveraged in each method. K-sigma uses only gross count rate information and had the poorest performance, which was to be expected in the presence of a rapidly changing background. mCEW, which constructs spectral regions of interest from its training data to perform detection, had greater sensitivity than k-sigma but still suffered when faced with the highly variable RADAI background. NSCRAD, too, constructs spectral windows to detect broad classes of anomalies, but it also incorporates information about its environment via a continually updating background model. This enables it to outperform mCEW and k-sigma, but it is susceptible to becoming stuck in a state of false alarm when abrupt background shifts occur. Conventional NMF achieved greater sensitivity by preserving the most spectral information compared to the other methods described and by making use of a broad library of templates of potential sources and a background model trained on many hours of data. The fact that Adaptive NMF was able to, in most test cases, match or improve upon the detection sensitivity of Conventional NMF is particularly encouraging due to the tradeoff in the information being utilized by each method. Adaptive NMF builds its background models from much smaller datasets — on the order of one hour’s worth of data — while Conventional NMF commonly uses many hours of data to build a complex model that is capable of capturing the richness of an environment over time in a single model. We have shown that in the case of highly variable background compositions, a flexible local model is not only sufficient but oftentimes superior to a highly complex static model.